OpenStudy (anonymous):

1. You roll a number cube once. Then you roll it again. What is the probability that you will get a 6 on the first roll and a number greater than 3 on the second roll? 1/12,1/6,2/3,1/9

Parth (parthkohli):

P(Getting a 6) = 1/6 P(Getting n>3) = 1/2 P(Getting both) = \(\Large \color{MidnightBlue}{\Rightarrow {1 \over 6} * {1 \over 2} = {1 \over 12} }\)

OpenStudy (anmolsingh):

the possible outcomes for such an event are given by 6^( no of dice or throws)=6^2=36 Now you have to use your brain. There will be only one case in which you wil get a 6 on the first dice and three cases where you will get a six on the first dice and 4,5,6 on the other =3events, so the probablity =3/36=1/12